in a reading ability test, with a sample of 100 cases, the mean score is 40 and

the standard deviation is 4.0. Assuming normality, What is the probability that a score will lie below score 46


1
Expert's answer
2022-06-14T13:12:46-0400

"P(X<46)=P(Z<\frac{46-40}{4})=P(Z<1.5)=0.9332"